# An Efficient Iterative Decoupling Method for Thermo-Poroelasticity Based on a Four-Field Formulation

### v2

@@ 2.1 A linear thermo-poroelasticity model
subject to the initial conditions:
$\bm{u}(\cdot,0)=\bm{u}^{0},\quad p(\cdot,0)=p^{0},\quad T(\cdot,0)=T^{0},$ (2.2)
and the homogeneous Dirichlet boundary conditions:
$\bm{u}=0,\quad p=0,\quad T=0,\quad\text{on }\partial\Omega\times J.$ (2.3)
@@
Note that in this paper considered coefficients and variables in the model are without units which is also known as dimensionless form.
@@
We assume the initial conditions satisfy $\bm{u}^{0}\in[H^{1}(\Omega)]^{d}$ and $p^{0},T^{0}\in L^{2}(\Omega)$, while the source terms satisfy $\bm{f}\in[L^{2}(\Omega)]^{d}$, $g\in L^{2}(\Omega)$, and $H_{s}\in L^{2}(\Omega)$. For simplicity, we further assume that $\bm{f}$, $g$, and $H_{s}$ are time-independent.
Define the following spaces
$\bm{V}=\bm{H}^{1}_{0}(\Omega),~{}Q=L^{2}(\Omega),~{}W=H_{0}^{1}(\Omega).$
@@
It is a well-established result that the following inf-sup condition [5] holds: there exists a constant $\zeta_{0}>0$ depending only on $\Omega$ such that
$\zeta_{0}\leq\inf_{\xi\in Q}\sup_{\bm{v}\in\bm{V}}\frac{|(\text{div }\bm{v},\xi)|}{\|\bm{v}\|_{H^{1}(\Omega)}\|\xi\|_{L^{2}(\Omega)}}.$ (2.5)
Multiplying with appropriate test functions, we obtain the weak form of the problem (2.4): For almost every $t\in(0,\tau]$, find $(\bm{u},~{}\xi,~{}p,~{}T)\in\bm{V}\times Q\times W\times W$ such that
@@
By the definition of $\xi=-\lambda\nabla\cdot\bm{u}+\alpha p+\beta T$, we have
@@ 2.2 A priori estimates of the weak solution
There exists a constant $C$ such that the Korn’s inequality
$C\|\bm{u}\|_{H^{1}(\Omega)}^{2}\leq\|\bm{\varepsilon}(\bm{u})\|_{L^{2}(\Omega)}^{2},\quad\forall\bm{u}\in[H^{1}_{0}(\Omega)]^{d},~{}d=2,3$
@@
$\|\xi(t)\|_{L^{2}(\Omega)}\leq C\big{(}2\mu\|\bm{\varepsilon}(u)\|_{L^{2}(\Omega)}+\|\bm{f}\|_{L^{2}(\Omega)}\big{)},$ (2.10)
where $C$ is a constant depending only on $\Omega$.
@@
$\zeta_{0}\|\xi\|_{L^{2}(\Omega)}=$ $\sup_{\bm{v}\in\bm{V}}\frac{|(\text{div }\bm{v},\xi)|}{\|\bm{v}\|_{H^{1}(\Omega)}}$
$\leq$ $\sup_{\bm{v}\in\bm{V}}\frac{|2\mu(\bm{\varepsilon}(\bm{u}),\bm{\varepsilon}(\bm{v}))+(\bm{f},\bm{v})|}{\|\bm{v}\|_{H^{1}(\Omega)}}$
$\leq$ $C_{K}\big{(}2\mu\|\bm{\varepsilon}(u)\|_{L^{2}(\Omega)}+\|\bm{f}\|_{L^{2}(\Omega)}\big{)},$
@@ 3 Numerical algorithms
This section introduces numerical algorithms derived from the four-field formulation previously discussed. Let $h$ denote the maximum diameter of all elements in the mesh, and for $h>0$, let $\mathcal{T}_{h}$ represent a family of triangulations of the domain $\Omega$ composed of triangular elements. The triangulations are assumed to be shape-regular and quasi-uniform. The discrete spaces for $\bm{u}$ and $\xi$ are denoted by $\bm{V}_{h}\times Q_{h}$, which are required to satisfy the following fundamental stability condition.
where $\gamma$ is independent of $h$, which means $V_{h}\times Q_{h}$ is a stable Stokes pair. As an example, we utilize the Taylor-Hood elements for the pair $(\bm{u},\xi)$, specifically $(\bm{P}_{2},P_{1})$ Lagrange finite elements, and $P_{1}$ Lagrange finite elements for both the fluid pressure $p$ and temperature $T$. The finite element spaces defined on $\mathcal{T}_{h}$ are as follows:
$\bm{V}_{h}=\left\{\bm{v}_{h}\in[H_{0}^{1}(\Omega)]^{2}\cap\bm{C}^{0}(\bar{\Omega}):\bm{v}_{h}|_{K}\in\mathbb{P}_{2}(K),\,\forall K\in\mathcal{T}_{h}\right\},$
@@
It is worth noting that other stable Stokes element pairs, such as the MINI element, could also be used.
@@ 5 Numerical experiments
This section presents numerical experiments to assess the computational accuracy and efficiency of the proposed algorithms in Section 3. The objective is to evaluate the performance of the two algorithms under different physical parameter settings. A uniform triangular partition, $\mathcal{T}_{h}$, is employed to generate the finite element mesh over the domain $\Omega=[0,1]\times[0,1]$, with the terminal time set to $\tau=0.01$. The time interval is defined as $J=[0,\tau]$. The experiments are conducted using the following benchmark problem. All computations are carried out using the FreeFEM++ software [16].
@@
$\bm{u}(x,y,t)=e^{-t}\left(\begin{array}[]{c}\sin(2\pi y)(\cos(2\pi x)-1)+\frac{1}{\mu+\lambda}\sin(\pi x)\sin(\pi y)\\ \sin(2\pi x)(1-\cos(2\pi y))+\frac{1}{\mu+\lambda}\sin(\pi x)\sin(\pi y)\\ \end{array}\right),$
@@
The corresponding boundary conditions and initial conditions are given as:
$\bm{u}(x,y,t)=\bm{0},~{}p(x,y,t)=0,~{}T(x,y,t)=0,\quad\mbox{on}~{}\partial\Omega,$
and
$\bm{u}(x,y,0)=\left(\begin{array}[]{c}\sin(2\pi y)(\cos(2\pi x)-1)+\frac{1}{\mu+\lambda}\sin(\pi x)\sin(\pi y)\\ \sin(2\pi x)(1-\cos(2\pi y))+\frac{1}{\mu+\lambda}\sin(\pi x)\sin(\pi y)\\ \end{array}\right),$
$p(x,y,0)=\sin(\pi x)\sin(\pi y),$
$T(x,y,0)=\sin(\pi x)\sin(\pi y).$
@@
For the coupled algorithm (Alg. 1), the time step size is set to $\Delta t=10^{-3}$, requiring 10 steps to reach the terminal time $\tau$. This means that the coupled system (3.2a)–(3.2d) is solved 10 times. For the decoupled algorithm (Alg. 2), two scenarios are considered: $\Delta t=5\times 10^{-3}$ and $\Delta t=10^{-2}$, corresponding to 2 and 1 time steps, respectively, to reach $\tau$. In these cases, the number of iterations is set to 5 and 10, respectively. Notably, the total number of system solves for (3.3)–(3.4) matches that of Alg. 1, ensuring an equivalent computational cost. This balance is achieved by appropriately selecting $\Delta t$ values and iteration counts. Despite this equivalence in system solves, the computational time for the decoupled algorithm is significantly reduced because solving smaller, independent subproblems is more efficient than solving the coupled system.
Three key observations emerge from the results: 1. Both the coupled and decoupled algorithms achieve optimal convergence rates in energy norm errors, as shown in Tables LABEL:tab:_K_Theta_0.1_nu_0.3 through LABEL:tab:a0=b0=c0=0. The decoupled algorithm matches the accuracy of the coupled algorithm while requiring less computational time. 2. The last entries in Tables LABEL:tab:_K_Theta_0.1_nu_0.3 through LABEL:tab:a0=b0=c0=0 reveal that increasing the number of iterations from 5 to 10 in the decoupled algorithm leads to slightly lower errors and higher convergence rates. This demonstrates that the algorithm’s accuracy improves with additional iterations. 3. Both algorithms maintain their effectiveness under extreme parameter values, as demonstrated in the subsequent subsections.
@@ 5.1 Tests for the parameter $\nu$
Table 1: Example 1: The errors and convergence orders with decreasing mesh sizes of coupled and iterative decoupling algorithms. $\nu=0.3$
@@
Alg.1, $\Delta t=10^{-3}$
$1/16$ 1.01028e-01 0.00 5.96583e-03 0.00 2.29827e-01 0.00 2.29827e-01 0.00
$1/32$ 2.54166e-02 1.99 1.47953e-03 2.01 1.09760e-01 1.07 1.09760e-01 1.07
$1/64$ 6.36417e-03 2.00 3.69057e-04 2.00 5.42034e-02 1.02 5.42034e-02 1.02
$1/128$ 1.59166e-03 2.00 9.20381e-05 2.00 2.70161e-02 1.00 2.70161e-02 1.00
Alg.2,$\Delta t=5\times 10^{-3}$,iters=5
$1/16$ 1.01030e-01 0.00 5.98023e-03 0.00 2.31732e-01 0.00 2.31732e-01 0.00
$1/32$ 2.54169e-02 1.99 1.48305e-03 2.01 1.10015e-01 1.07 1.10015e-01 1.07
$1/64$ 6.36426e-03 2.00 3.69788e-04 2.00 5.42370e-02 1.02 5.42370e-02 1.02
$1/128$ 1.59168e-03 2.00 9.21223e-05 2.01 2.70211e-02 1.01 2.70211e-02 1.01
Alg.2,$\Delta t=10^{-2}$,iters=10
$1/16$ 1.01031e-01 0.00 5.99684e-03 0.00 2.34369e-01 0.00 2.34369e-01 0.00
$1/32$ 2.54172e-02 1.99 1.48591e-03 2.01 1.10312e-01 1.09 1.10312e-01 1.09
$1/64$ 6.36419e-03 2.00 3.69110e-04 2.01 5.42412e-02 1.02 5.42412e-02 1.02
$1/128$ 1.59156e-03 2.00 9.07679e-05 2.02 2.70075e-02 1.01 2.70075e-02 1.01
Table 2: Example 1: The errors and convergence orders with decreasing mesh sizes of coupled and iterative decoupling algorithms, Alg.1 with $\Delta t=t=5\times 10^{-3}$ and $\Delta t=10^{-2}$, $\nu=0.3$
Alg.1,$\Delta t=5\times 10^{-3}$
$1/16$ 1.01125e-01 0.00 6.84677e-03 0.00 2.31369e-01 0.00 2.31369e-01 0.00
$1/32$ 2.54287e-02 1.99 1.59617e-03 2.10 1.09959e-01 1.07 1.09959e-01 1.07
$1/64$ 6.36566e-03 2.00 3.83650e-04 2.06 5.42156e-02 1.02 5.42156e-02 1.02
$1/128$ 1.59179e-03 2.00 9.32580e-05 2.04 2.70112e-02 1.01 2.70112e-02 1.01
Alg.1,$\Delta t=10^{-2}$
$1/16$ 1.01127e-01 0.00 6.85872e-03 0.00 2.33940e-01 0.00 2.33940e-01 0.00
$1/32$ 2.54290e-02 1.99 1.59902e-03 2.10 1.10284e-01 1.08 1.10284e-01 1.08
$1/64$ 6.36567e-03 2.00 3.83660e-04 2.06 5.42396e-02 1.02 5.42396e-02 1.02
$1/128$ 1.59175e-03 2.00 9.26565e-05 2.05 2.70074e-02 1.01 2.70074e-02 1.01
Table 3: Example 1: The errors and convergence orders with decreasing mesh sizes of coupled and iterative decoupling algorithms. $\nu=0.49999$
Alg.1, $\Delta t=10^{-3}$
$1/16$ 1.00295e-01 0.00 9.63830e-03 0.00 2.15490e-01 0.00 2.15490e-01 0.00
$1/32$ 2.52279e-02 1.99 2.38247e-03 2.02 1.07906e-01 1.00 1.07906e-01 1.00
$1/64$ 6.31671e-03 2.00 5.94250e-04 2.00 5.39732e-02 1.00 5.39732e-02 1.00
$1/128$ 1.57979e-03 2.00 1.48471e-04 2.00 2.69891e-02 1.00 2.69891e-02 1.00
Alg.2, $\Delta t=5\times 10^{-3}$, iters=5
$1/16$ 1.00295e-01 0.00 9.63812e-03 0.00 2.15492e-01 0.00 2.15492e-01 0.00
$1/32$ 2.52279e-02 1.99 2.38236e-03 2.02 1.07907e-01 1.00 1.07907e-01 1.00
$1/64$ 6.31671e-03 2.00 5.94164e-04 2.00 5.39735e-02 1.00 5.39735e-02 1.00
$1/128$ 1.57979e-03 2.00 1.48397e-04 2.00 2.69893e-02 1.00 2.69893e-02 1.00
Alg.2, $\Delta t=10^{-2}$, iters=10
$1/16$ 1.00295e-01 0.00 9.63791e-03 0.00 2.15495e-01 0.00 2.15495e-01 0.00
$1/32$ 2.52279e-02 1.99 2.38224e-03 2.02 1.07907e-01 1.00 1.07907e-01 1.00
$1/64$ 6.31671e-03 2.00 5.94068e-04 2.00 5.39738e-02 1.00 5.39738e-02 1.00
$1/128$ 1.57979e-03 2.00 1.48328e-04 2.00 2.69895e-02 1.00 2.69895e-02 1.00
@@
In this subsection, we evaluate the performance of the algorithms under various settings of the Poisson ratio $\nu$. For the other parameters, we fix $E=1$, $c_{0}=a_{0}=0.2$, $b_{0}=0.1$, $\alpha=\beta=0.1$, and $\bm{K}=\bm{\Theta}=0.1\bm{I}$, where $\bm{I}$ denotes the $2\times 2$ identity matrix. The data shown in Table LABEL:tab:_K_Theta_0.1_nu_0.3, along with the subsequent tables LABEL:tab:_K_1e-6-LABEL:tab:a0=b0=c0=0, are based on a setting where $\nu=0.3$. These tables represent the situation where the thermo-poroelastic material is compressible.
In Table LABEL:tab:_Alg.1_Dt_0.005_0.01, we present the performance of Algorithm 1 at time steps $\Delta t=5\times 10^{-3}$ and $\Delta t=1\times 10^{-2}$. When compared with Algorithm 2 at the same time steps in Table LABEL:tab:_K_Theta_0.1_nu_0.3, the errors of Algorithm 1 are slightly larger than or comparable to those of Algorithm 2.
@@
In Table LABEL:tab:_nu_0.49999, we set a constant Poisson ratio of $\nu=0.49999$, while keeping all other physical parameters fixed. Since the Poisson ratio $\nu$ is close to $0.5$, which in other words means $\lambda\to\infty$, the thermo-poroelastic material behaves nearly incompressibly, leading the mixed linear elasticity model to approach the incompressible Stokes model. The results presented in Table LABEL:tab:_nu_0.49999 demonstrate that the energy-norm errors for all algorithms exhibit optimal orders as the material approaches incompressibility, thereby emphasizing the robustness of these algorithms with respect to the parameter $\nu$.

### v3

@@ 2.1 A linear thermo-poroelasticity model
Proper boundary and initial conditions should be provided in order to ensure the existence and uniqueness of the solution. In this paper, we consider a mixed partial Neumann and partial Dirichlet conditions for $\bm{u}$ and the Dirichlet conditions for $p$ and $T$. Assume $\partial\Omega=\Gamma_{d}+\Gamma_{n}$ with $\Gamma_{d}\neq\partial\Omega$. Here, $\Gamma_{d}$ and $\Gamma_{n}$ denote the Dirichlet boundary and Neumann boundary for $\bm{u}$, respectively. Without loss of generality, the boundary conditions are assumed to be homogeneous. Specifically, the boundary conditions are given by
$\bm{u}=0,$ on $\text{on }\Gamma_{d},$ (2.2)
$(2\mu\bm{\varepsilon}(\bm{u})+\lambda\nabla\cdot\bm{u}\bm{I}-\alpha p\bm{I}-\beta T\bm{I})\bm{n}=0,$ on $\text{on }\Gamma_{n},$
$p=0,$ on $\text{on }\partial\Omega,$
$T=0,$ on $\text{on }\partial\Omega.$
Here, $\bm{n}$ is the unit outward normal to the boundary. The initial conditions are given by
$\bm{u}(\cdot,0)=\bm{u}^{0},\quad p(\cdot,0)=p^{0},\quad T(\cdot,0)=T^{0}.$ (2.3)
@@
Note that in this paper, the considered coefficients and variables in the model are without units, which is also known as the dimensionless form.
@@
We assume the initial conditions satisfy $\bm{u}^{0}\in[H^{1}(\Omega)]^{d}$ and $p^{0},T^{0}\in L^{2}(\Omega)$, while the source terms satisfy $\bm{f}\in[L^{2}(\Omega)]^{d}$, $g\in L^{2}(\Omega)$, and $H_{s}\in L^{2}(\Omega)$. For simplicity, we further assume that $\bm{f}$, $g$, and $H_{s}$ are time-independent. Define the following spaces
$\bm{V}=\bm{H}^{1}_{0,\Gamma_{d}}(\Omega),~{}Q=L^{2}(\Omega),~{}W=H_{0}^{1}(\Omega).$
@@
$\zeta_{0}\leq\inf_{\xi\in L^{2}(\Omega)}\sup_{\bm{v}\in\bm{V}}\frac{|(\text{div }\bm{v},\xi)|}{\|\bm{v}\|_{H^{1}(\Omega)}\|\xi\|_{L^{2}(\Omega)}}.$ (2.5)
Its proof can be found in Quarteroni’s book [22] or Cai’s thesis [10] (Lemma 2.2). Given $\tau>0$, a 4-tuple $(\bm{u},\xi,p,T)$ with
@@
is called a weak solution of problem (2.4), if for almost every $t\in(0,\tau]$, and $\forall(\bm{v},~{}\phi,~{}q,~{}S)\in\bm{V}\times Q\times W\times W$, there holds
@@
Here, these equations can be derived by multiplying the differential equations with appropriate test functions. By the definition of $\xi=-\lambda\nabla\cdot\bm{u}+\alpha p+\beta T$, we have
@@ 2.2 A priori estimates of the weak solution
There exists a constant $C$ such that Korn’s inequality
$C\|\bm{u}\|_{H^{1}(\Omega)}^{2}\leq\|\bm{\varepsilon}(\bm{u})\|_{L^{2}(\Omega)}^{2},\quad\forall\bm{u}\in[H^{1}_{0,\Gamma_{d}}(\Omega)]^{d},~{}d=2,3$
@@
$\|\xi(t)\|_{L^{2}(\Omega)}\leq C\big{(}2\mu\|\bm{\varepsilon}(\bm{u})\|_{L^{2}(\Omega)}+\|\bm{f}\|_{L^{2}(\Omega)}\big{)},$ (2.10)
where $C$ is a constant depending only on $\Omega$ and $\Gamma_{d}$.
@@
$\zeta_{0}\|\xi\|_{L^{2}(\Omega)}\leq$ $\sup_{\bm{v}\in\bm{V}}\frac{|(\text{div }\bm{v},\xi)|}{\|\bm{v}\|_{H^{1}(\Omega)}}$
$\leq$ $\sup_{\bm{v}\in\bm{V}}\frac{|2\mu(\bm{\varepsilon}(\bm{u}),\bm{\varepsilon}(\bm{v}))|+|(\bm{f},\bm{v})|}{\|\bm{v}\|_{H^{1}(\Omega)}}$
$\leq$ $C_{K}\big{(}2\mu\|\bm{\varepsilon}(\bm{u})\|_{L^{2}(\Omega)}+\|\bm{f}\|_{L^{2}(\Omega)}\big{)},$
@@ 3 Numerical algorithms
This section introduces numerical algorithms derived from the four-field formulation previously discussed. Let $h$ denote the maximum diameter of all elements in the mesh, and for $h>0$, let $\mathcal{T}_{h}$ represent a family of triangulations of the domain $\Omega$ composed of triangular elements. The triangulations are assumed to be shape-regular and quasi-uniform. The finite element spaces defined on $\mathcal{T}_{h}$ are as follows:
$\bm{V}_{h}=\left\{\bm{v}_{h}\in[H_{0,\Gamma_{d}}^{1}(\Omega)]^{2}\cap\bm{C}^{0}(\bar{\Omega}):\bm{v}_{h}|_{K}\in\mathbb{P}_{2}(K),\,\forall K\in\mathcal{T}_{h}\right\},$
@@
The discrete spaces for $\bm{u}$ and $\xi$ are denoted by $\bm{V}_{h}\times Q_{h}$, which are required to satisfy the following fundamental stability condition.
where $\gamma$ is independent of $h$, which means $V_{h}\times Q_{h}$ is a stable Stokes pair. As an example, we utilize the Taylor-Hood elements for the pair $(\bm{u},\xi)$, specifically $(\bm{P}_{2},P_{1})$ Lagrange finite elements, and $P_{1}$ Lagrange finite elements for both the fluid pressure $p$ and temperature $T$. It is worth noting that other stable Stokes element pairs, such as the MINI element, could also be used.
@@ 5 Numerical experiments
This section presents numerical experiments to assess the computational accuracy and efficiency of the proposed algorithms in Section 3. The objective is to evaluate the performance of the two algorithms under different physical parameter settings. A uniform triangular partition, $\mathcal{T}_{h}$, is employed to generate the finite element mesh over the domain $\Omega=[0,1]\times[0,1]$, with the terminal time set to $\tau=0.01$ The time interval is defined as $J=[0,\tau]$. The experiments are conducted using the following benchmark problem. All computations are carried out using the FreeFEM++ software [16].
@@
$\bm{u}(x,y,t)=e^{-t}\left(\begin{array}[]{c}\sin(2\pi y)(\cos(2\pi x)-1)+\frac{1}{\mu+\lambda}\sin(\pi x)\sin(\pi y)\\ \sin(2\pi x)(1-\cos(2\pi y))+\frac{1}{\mu+\lambda}\sin(\pi x)\sin(\pi y)\\ \end{array}\right),$ (5.1)
@@
We impose Dirichlet boundary conditions $\bm{u}=\bm{0}$ on the boundary segment $\Gamma_{d}=\{(x,y)\,|\,0\leq y\leq 1,\,x=0\text{ or }1\}$, and Neumann boundary conditions on $\Gamma_{n}=\{(x,y)\,|\,0\leq x\leq 1,\,y=0\text{ or }1\}$. For the pressure $p$ and temperature $T$, homogeneous Dirichlet boundary conditions are prescribed on the entire boundary $\partial\Omega$. The initial conditions for $\bm{u}^{0}$, $p^{0}$, and $T^{0}$ at time $t=0$ are given by the expression in (5.1).
@@
For the coupled algorithm (Alg. 1), the time step size is set to $\Delta t=10^{-3}$, requiring 10 steps to reach the terminal time $\tau$. This means that the coupled system (3.2a)–(3.2d) is solved 10 times. For the decoupled algorithm (Alg. 2), two scenarios are considered: $\Delta t=5\times 10^{-3}$ and $\Delta t=10^{-2}$, corresponding to 200 and 100 time steps, respectively, to reach $\tau$. In these cases, the number of iterations is set to 5 and 10, respectively. Notably, the total number of system solves for (3.3)–(3.4) matches that of Alg. 1, ensuring an equivalent computational cost. This balance is achieved by appropriately selecting $\Delta t$ values and iteration counts. Despite this equivalence in system solves, the computational time for the decoupled algorithm is significantly reduced because solving smaller, independent subproblems is more efficient than solving the coupled system.
From the results, we have the following three key observations: 1. Both the coupled and decoupled algorithms attain optimal convergence rates in the energy norm, as demonstrated in Tables LABEL:tab:_K_Theta_0.1_nu_0.3 to LABEL:tab:a0=b0=c0=0. Notably, the decoupled algorithm achieves comparable accuracy to the coupled counterpart while reducing computational cost. For instance, CPU times (in seconds) for each algorithm are presented in Table LABEL:tab:_K_Theta_0.1_nu_0.3 for comparison. 2. The final entries in Tables LABEL:tab:_K_Theta_0.1_nu_0.3 through LABEL:tab:a0=b0=c0=0 indicate that increasing the number of iterations in the decoupled algorithm from 5 to 10 results in slightly improved accuracy and enhanced convergence rates. This highlights the potential for accuracy gains with more iterations in each time step. 3. Both algorithms exhibit robust performance even under extreme parameter regimes, as evidenced in the subsequent numerical experiments.
@@ 5.1 Tests for the parameter $\nu$
Table 1: The errors and convergence orders with decreasing mesh sizes. $\nu=0.3$.
$h$ $\|\bm{u}\|_{H^{1}}$ orders $\|\xi\|_{L^{2}}$ orders $\|p\|_{H^{1}}$ orders $\|T\|_{H^{1}}$ orders CPU time(s)
Alg.1,
$\Delta t=10^{-3}$
$1/16$ 1.00607e-01 0.00 6.00958e-03 0.00 2.28033e-01 0.00 2.28033e-01 0.00 3.32
$1/32$ 2.53649e-02 1.99 1.48475e-03 2.02 1.09515e-01 1.06 1.09515e-01 1.06 8.87
$1/64$ 6.35806e-03 2.00 3.69965e-04 2.00 5.41728e-02 1.02 5.41728e-02 1.02 31.26
$1/128$ 1.59098e-03 2.00 9.22429e-05 2.00 2.70126e-02 1.00 2.70126e-02 1.00 153.86
Alg.2,
$\Delta t=5\times 10^{-3}$,
iters=5
$1/16$ 1.00607e-01 0.00 6.01533e-03 0.00 2.28993e-01 0.00 2.28993e-01 0.00 2.03
$1/32$ 2.53650e-02 1.99 1.48612e-03 2.02 1.09638e-01 1.06 1.09638e-01 1.06 6.67
$1/64$ 6.35808e-03 2.00 3.70229e-04 2.01 5.41866e-02 1.02 5.41866e-02 1.02 19.26
$1/128$ 1.59097e-03 2.00 9.22425e-05 2.00 2.70136e-02 1.00 2.70136e-02 1.00 79.37
Alg.2,
$\Delta t=10^{-2}$,
iters=10
$1/16$ 1.00608e-01 0.00 6.02199e-03 0.00 2.30330e-01 0.00 2.30330e-01 0.00 2.27
$1/32$ 2.53650e-02 1.99 1.48635e-03 2.02 1.09772e-01 1.07 1.09772e-01 1.07 7.49
$1/64$ 6.35790e-03 2.00 3.68857e-04 2.01 5.41797e-02 1.02 5.41797e-02 1.02 21.37
$1/128$ 1.59078e-03 2.00 9.07180e-05 2.02 2.70029e-02 1.00 2.70029e-02 1.00 85.17
Table 2: The errors and convergence orders with decreasing mesh sizes. $\nu=0.49999$.
@@
Alg.1,
$\Delta t=10^{-3}$
$1/16$ 9.99038e-02 0.00 9.71217e-03 0.00 2.15222e-01 0.00 2.15222e-01 0.00
$1/32$ 2.51776e-02 1.99 2.38401e-03 2.03 1.07872e-01 1.00 1.07872e-01 1.00
$1/64$ 6.31033e-03 2.00 5.93579e-04 2.01 5.39689e-02 1.00 5.39689e-02 1.00
$1/128$ 1.57899e-03 2.00 1.48251e-04 2.00 2.69886e-02 1.00 2.69886e-02 1.00
Alg.2,
$\Delta t=5\times 10^{-3}$,
iters=5
$1/16$ 9.99038e-02 0.00 9.71217e-03 0.00 2.15234e-01 0.00 2.15234e-01 0.00
$1/32$ 2.51776e-02 1.99 2.38401e-03 2.03 1.07874e-01 1.00 1.07874e-01 1.00
$1/64$ 6.31033e-03 2.00 5.93579e-04 2.01 5.39692e-02 1.00 5.39692e-02 1.00
$1/128$ 1.57899e-03 2.00 1.48251e-04 2.00 2.69886e-02 1.00 2.69886e-02 1.00
Alg.2,
$\Delta t=10^{-2}$,
iters=10
$1/16$ 9.99038e-02 0.00 9.71217e-03 0.00 2.15248e-01 0.00 2.15248e-01 0.00
$1/32$ 2.51776e-02 1.99 2.38401e-03 2.03 1.07876e-01 1.00 1.07876e-01 1.00
$1/64$ 6.31033e-03 2.00 5.93579e-04 2.01 5.39694e-02 1.00 5.39694e-02 1.00
$1/128$ 1.57899e-03 2.00 1.48251e-04 2.00 2.69888e-02 1.00 2.69888e-02 1.00
@@
In this subsection, we evaluate the performance of the algorithms under various settings of the Poisson ratio $\nu$. For the other parameters, we fix $E=1$, $c_{0}=a_{0}=0.2$, $b_{0}=0.1$, $\alpha=\beta=0.1$, and $\bm{K}=\bm{\Theta}=0.1\bm{I}$, where $\bm{I}$ denotes the $2\times 2$ identity matrix. The data shown in Table LABEL:tab:_K_Theta_0.1_nu_0.3, along with the subsequent tables LABEL:tab:_K_1e-6-LABEL:tab:a0=b0=c0=0, are based on a setting where $\nu=0.3$. These tables represent the situation where the thermo-poroelastic material is compressible. We also report the CPU times for the different algorithms. In particular, comparison between the coupled and decoupled algorithms reveals that the decoupled approaches incur lower computational costs, primarily because each subproblem in the decoupled scheme involves significantly smaller system sizes compared to the fully coupled formulation.
@@
In Table LABEL:tab:_nu_0.49999, we set a constant Poisson ratio of $\nu=0.49999$, while keeping all other physical parameters fixed. Since the Poisson ratio $\nu$ is close to $0.5$ which in other words means $\lambda\to\infty$, the thermo-poroelastic material behaves nearly incompressibly, leading the mixed linear elasticity model to approach the incompressible Stokes model. The results presented in Table LABEL:tab:_nu_0.49999 demonstrate that the energy-norm errors for all algorithms exhibit optimal orders as the material approaches incompressibility, thereby emphasizing the robustness of these algorithms with respect to the parameter $\nu$.